a) A thin metal plate bounded by the π‘₯-axis and the lines π‘₯ = 0 and π‘₯ = 1 and


stretching to infinity in the 𝑦-direction has its upper and lower faces perfectly


insulated and its vertical edges and edge at infinity are maintained at the constant


temperature 0 °𝐢, while over the base temperature of 50 °𝐢 is maintained. Find


the steady state temperature 𝑒(π‘₯,𝑑).


[5]


b) If 𝑓(π‘₯𝑦^2, 𝑧 βˆ’ 2π‘₯) = 0 then prove that π‘₯ 𝑧x βˆ’1/2𝑦 𝑧𝑦 = 2 π‘₯

1
Expert's answer
2022-01-05T14:02:51-0500

b)

prove:

"x\frac{\partial z}{\partial x}-\frac{\partial z}{\partial y}\frac{y}{2}=2x"


Differentiating "\ud835\udc53(\ud835\udc65\ud835\udc66^2, \ud835\udc67 \u2212 2\ud835\udc65) = 0" , we get:

"y^2f_1+(\frac{\partial z}{\partial x}-2)f_2=0"


"2xyf_1+\frac{\partial z}{\partial y}f_2=0"


where f1 and f2 are partial derivatives of f with respect to its first input and second input. When we multiply the first equality above by 2x and the second by y:

"2x(\frac{\partial z}{\partial x}-2)=y\frac{\partial z}{\partial y}\implies x\frac{\partial z}{\partial x}-\frac{\partial z}{\partial y}\frac{y}{2}=2x"


a)

solution of two dimensional heat equation "\frac{\partial^2 u}{\partial x^2}+\frac{\partial^2 u}{\partial y^2}=0" is


"u(x,y)=(Acos\lambda x+Bsin \lambda x)(Ce^{\lambda y}+De^{-\lambda y})"

boundary conditions:

"u(x,\infin)=0"

"u(x,0)=50"

"u(0,y)=0"

"u(1,y)=0"

"0\le x\le 1,0\le y< \infin"


then:

"u(0,y)=A(Ce^{\lambda y}+De^{-\lambda y})=0\implies A=0"

"u(x,y)=Bsin \lambda x(Ce^{\lambda y}+De^{-\lambda y})"


"u(1,y)=Bsin \lambda (Ce^{\lambda y}+De^{-\lambda y})=0\implies \lambda=n\pi"

then:

"u(x,y)=Bsin (\pi nx)(Ce^{\pi n y}+De^{-\pi n y})"

"u(x,\infin)=Bsin (\pi nx)Ce^{\pi n y}=0\implies C=0"


"u(x,y)=B_1sin (\pi nx)e^{-\pi n y}"

where B1 = BD


the most general solution:

"u(x,y)=\sum B_nsin (\pi nx)e^{-\pi n y}"

where

"B_n=\int^1_0 u(x,0)sin (\pi nx)=50\int^1_0 sin (\pi nx)="


"=-\frac{50}{\pi n}cos(n\pi x)|^1_0=\frac{50}{\pi n}(1-cos(n\pi))"


for odd n:

"u(x,y)=\sum\frac{100}{\pi n}sin (\pi nx)e^{-\pi n y}"


for even n:

"u(x,y)=0"